OpenStudy (anonymous):

Help plz 4 medal and fan

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

x^-40 * x^2 = x^-40+2 =x^38

OpenStudy (hoblos):

first \[(x^a)^b = x^{a \times b}\] so \[(x^5)^{-8} = ?\]

OpenStudy (anonymous):

how can you times X @hoblos

OpenStudy (hoblos):

you mean if we have \[x^a x^b\] ?

OpenStudy (anonymous):

no (X^5)^-8 how do u times x?

OpenStudy (anonymous):

urggg i hate this

OpenStudy (hoblos):

the general formula says \[(x^a)^b = x^{ab}\] so here \[(x^5)^{-8} = x^{5(-8)} = x^{-40}\]

OpenStudy (hoblos):

now \[x^{-40} x^2 = ?\]

OpenStudy (anonymous):

umm donno

OpenStudy (anonymous):

sry im freakin stupid

OpenStudy (hoblos):

\[x^a x^b = x^{a + b}\] so here \[x^{-40} x^2 = x^{-40+2} = ?\]

OpenStudy (anonymous):

x^-42

OpenStudy (hoblos):

don't say that you are NOT just try and you will be able to solve it

OpenStudy (hoblos):

-40+2 = -38

OpenStudy (anonymous):

ive failed 7 years of mathim in 9th grade

OpenStudy (anonymous):

There's a formula that if a term is raised to the power of something outside brackets, the power outside brackets will be multiplied. Even if the power inside is 1 it'll get multiplied Here \[(x ^{5})^{-8}= x ^{5\times-8}=x ^{-40}\] rest has been explained by hoblos

OpenStudy (hoblos):

that's mot a shame , Bill Gates failed to so must not stop trying

OpenStudy (the_fizicx99):

He did :O

OpenStudy (the_fizicx99):

\[x ^{-b} = \frac{ 1 }{ x^b }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

whats that

OpenStudy (the_fizicx99):

Another rule, it's how you make the negative exponent positive

OpenStudy (anonymous):

oh gosh dats the last thing i want can ya give me the answer plz i would like to move on

OpenStudy (hoblos):

you will use that after a bit @luvalpaca now \[x^{-40+2} = ?\]

OpenStudy (anonymous):

grrrr

OpenStudy (anonymous):

just give me the answer plz

OpenStudy (hoblos):

just try once more -40+2 =-38 so \[x^{-40+2} = ?\]

OpenStudy (anonymous):

ppl these days

OpenStudy (anonymous):

ive study my ____ off on khan academy and i still dont get what the heck to do just give me the answer im gettin mad plz

OpenStudy (the_fizicx99):

Answer what hoblos asked you and you'll get your answer.

OpenStudy (anonymous):

I DONT KNOW WHAT THE flutter IT ISSSSSSSSSSSSSSSSSSSSSS

OpenStudy (anonymous):

Exactly.. I've made a mistake in my first answer. See if you can correct that luvalpaca

OpenStudy (the_fizicx99):

-40 + 2 = ?

OpenStudy (hoblos):

calm down @luvalpaca it just need one second for you to get it, just try it once more and tell me what you get

OpenStudy (anonymous):

im good bye

OpenStudy (hoblos):

wait a second

OpenStudy (hoblos):